# WW-FL: Secure and Private Large-Scale Federated Learning

### v3

@@ 3.2. MPC Functionalities in WW-FL
The MPC functionalities utilized in WW-FL are listed in Tab. 3 and discussed in the following. While Share and Reshare are used for generating the secret shares as per the underlying MPC semantics, Reveal is used to reconstruct the secret towards a designated party. The Train and Predict functionalities correspond to PPML training and inference protocols, respectively. Similarly, Agg denotes the secure aggregation functionality in FL (cf. §2). In WW-FL, these functionalities are realized using Meta’s CrypTen framework, in which two semi-honest MPC servers carry out the computation with the help of a trusted third party that provides correlated randomness (Damgård et al., 2012; Demmler et al., 2015; Riazi et al., 2018; Knott et al., 2021). However, WW-FL is not bound to any specific MPC setting and each of the MPC clusters as well as the global servers could be instantiated independently using any MPC protocol, as detailed further in §3.3 and §3.5.
Table 3. MPC functionalities used in WW-FL framework.
Algorithm Input(s) Description
Share ${D},{V}$ Generates secret shares of ${D}$ as per target ${V}$’s sharing semantics (${\langle{D}\rangle}_{{V}}$).
@@
Reshare ${\langle{D}\rangle}_{{U}},{V}$ Converts secret shares of ${\langle{D}\rangle}$ from source ${U}$’s sharing semantics to target ${V}$.
Train ${\langle{W}\rangle},{\langle{D}\rangle}$ Performs PPML Training on ML model ${W}$ using the secret-shared data ${\langle{D}\rangle}$.
Reveal ${\langle{D}\rangle}_{{U}},{V}$ Reconstructs secret-shared data ${\langle{D}\rangle}$ towards members in target set ${V}$.
3.3. Threat Model and Privacy Analysis
Given the hybrid nature of our framework, we consider a mixed adversarial corruption in WW-FL and assume a centralized adversary $\mathcal{A}$ orchestrating the corruption [^4] This is a stronger notion than having multiple independent adversaries., which allows the nature of the corruption to differ amongst MPC clusters including the global servers. We model this by defining a corruption strategy $\mathsf{St}_{\gamma}$ for the entities in the set $\gamma\in\{\mathcal{G},\mathcal{M}_{i},\mathcal{C}\}_{i\in\mathsf{N}_{\footnotesize\mathcal{M}}}$. For example, $\mathcal{A}$ can choose to maliciously corrupt certain MPC clusters $\mathcal{M}_{i}$ (reflecting an unjustified trust assumption in the cluster), while corrupting the rest, including $\mathcal{G}$, in a semi-honest manner. $\mathcal{A}$ also corrupts a minority or majority of entities in a set, depending on the number of trustworthy entities (assuming at least one trustworthy entity in each set). The security spectrum of WW-FL is summarized in Tab. 4, ranging from weaker semi-honest cases to the strongest all-malicious setting based on the nature of entity corruption.
@@ 3.3. Threat Model and Privacy Analysis
Table 4. Spectrum of security achievable in  WW-FL based on the corruption strategy of the adversary. Here $0<t<\mathsf{N}_{\mathcal{M}}$.
@@
I-A $\mathsf{semi\mbox{-}honest}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{semi\mbox{-}honest}$ $\mathsf{semi\mbox{-}honest}$
I-B $\mathsf{semi\mbox{-}honest}$ $t$ $\mathsf{malicious}$ $\mathsf{semi\mbox{-}honest}$
I-C $\mathsf{semi\mbox{-}honest}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{malicious}$ $\mathsf{semi\mbox{-}honest}$
II-A $\mathsf{malicious}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{semi\mbox{-}honest}$ $\mathsf{semi\mbox{-}honest}$
II-B $\mathsf{malicious}$ $t$ $\mathsf{malicious}$ $\mathsf{semi\mbox{-}honest}$
II-C $\mathsf{malicious}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{malicious}$ $\mathsf{semi\mbox{-}honest}$
III-A $\mathsf{semi\mbox{-}honest}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{semi\mbox{-}honest}$ $\mathsf{malicious}$
III-B $\mathsf{semi\mbox{-}honest}$ $t$ $\mathsf{malicious}$ $\mathsf{malicious}$
III-C $\mathsf{semi\mbox{-}honest}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{malicious}$ $\mathsf{malicious}$
IV-A $\mathsf{malicious}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{semi\mbox{-}honest}$ $\mathsf{malicious}$
IV-B $\mathsf{malicious}$ $t$ $\mathsf{malicious}$ $\mathsf{malicious}$
IV-C $\mathsf{malicious}$ $\mathsf{N}_{\mathcal{M}}$ $\mathsf{malicious}$ $\mathsf{malicious}$
@@
We model the FL training process in WW-FL as an ideal functionality $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ which is depicted in Alg. 2. For simplicity, we assume that the clients have secret-shared their training data with their respective MPC clusters via the respective Share protocol. In each iteration $t$, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ receives the shares of the current model $W_{t-1}$ from the global servers ($\mathcal{G}$) and the shares of the training data from each cluster’s servers in $\mathcal{M}_{i}$ for $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$. Using these shares, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ reconstructs the current global model and the training data for each cluster. Next, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ trains the model $W_{t-1}$ using the data from each cluster, resulting in a cluster model. These cluster models are then combined using an aggregation algorithm to obtain the updated global model $W_{t}$ for this iteration. Finally, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ generates the secret-shares of $W_{t}$ among the global servers ($\mathcal{G}$) and the cluster servers ($\mathcal{M}_{i}$) using their respective sharing semantics.
@@
0:  $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$, $\mathcal{G},\mathcal{M}$ $\mathcal{M}=\bigcup_{i}\mathcal{M}_{i}$; $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$
@@
[…]
@@
3:     data reconstruction: $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ receives ${\langle{D}_{t}^{\mathcal{M}_{i}}\rangle}$ from $\mathcal{M}_{i}$ and reconstructs the cluster data ${D}_{t}^{\mathcal{M}_{i}}$.${D}_{t}^{\mathcal{M}_{i}}$ corresponds to collective data at $\mathcal{M}_{i}$
@@
6:  aggregation: $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ aggregates $\{W_{t}^{\mathcal{M}_{i}}\}_{i\in\mathsf{N}_{\footnotesize\mathcal{M}}}$ using the respective secure aggregation algorithm to obtain $W_{t}{}$.
@@
8:  secret-sharing (cluster servers): $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ generates ${\langle W_{t}\rangle}_{\mathcal{M}_{i}}$ from $W_{t}{}$ for $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$.
The security of WW-FL is stated in Theorem 3.1, and we give detailed security proof in §A.2.
@@
Algorithm FL-Train (Alg. 1) securely realises the $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ functionality (Alg. 2) in the $\mathcal{F}_{\textsf{MPC}}$-hybrid model, assuming the presence of a static adversary $\mathcal{A}$, who corrupts the entities in $\gamma\in\{\mathcal{G},\mathcal{M}_{i},\mathcal{C}\}_{i\in\mathsf{N}_{\footnotesize\mathcal{M}}}$ using a predefined corruption strategy $\mathsf{St}_{\gamma}$.
3.4. Private Inference in WW-FL

### v4

@@ 3.2 MPC Functionalities in WW-FL
The MPC functionalities utilized in WW-FL are listed in Tab. 3 and discussed in the following. While Share and Reshare are used for generating the secret shares as per the underlying MPC semantics, Reveal is used to reconstruct the secret towards a designated party. The Train and Predict functionalities correspond to PPML training and inference protocols, respectively. Similarly, Agg denotes the secure aggregation functionality in FL (cf. §2). In WW-FL, these functionalities are realized using Meta’s CrypTen framework, in which two semi-honest MPC servers carry out the computation with the help of a trusted third party that provides correlated randomness [DPSZ12, DSZ15, RWT+18, KVH+21]. However, WW-FL is not bound to any specific MPC setting and each of the MPC clusters as well as the global servers could be instantiated independently using any MPC protocol, as detailed further in §3.3 and §3.5.
$\mathcal{F}_{\textnormal{{Reshare}}}$ Ideal Functionality
Given the hybrid nature of WW-FL, the security of the framework heavily relies on the correct instantiation of the ideal functionality for the Reshare algorithm, denoted as $\mathcal{F}_{\textnormal{{Reshare}}}$. This ideal functionality is presented in Alg. 2.
Algorithm 2 $\mathcal{F}_{\textnormal{{Reshare}}}$: Ideal Functionality for Reshare
0:  $\mathcal{P}:{\langle{D}\rangle}_{{U}}$, $\mathcal{P}^{\prime}:\bot$
0:  $\mathcal{P}:\bot$, $\mathcal{P}^{\prime}:{\langle{D}\rangle}_{{V}}$
@@
1:  secret reconstruction: $\mathcal{F}_{\textnormal{{Reshare}}}$ receives ${\langle{D}\rangle}_{{U}}$ from $\mathcal{P}$ and reconstructs secret D{D} using the sharing semantics of source U{U}.
2:  share generation: $\mathcal{F}_{\textnormal{{Reshare}}}$ computes ${\langle{D}\rangle}_{{V}}$ from D{D} using sharing semantics of target V{V}.
3:  share distribution: $\mathcal{F}_{\textnormal{{Reshare}}}$ sends the secret-shares of ${\langle{D}\rangle}_{{V}}$ to parties in $\mathcal{P}^{\prime}$ in accordance with the sharing semantics of target V{V}.
The purpose of $\mathcal{F}_{\textnormal{{Reshare}}}$ is to securely convert secret shares of data D{D} from the sharing semantics of the source U{U}, defined over a party set $\mathcal{P}$ (i.e., ${\langle{D}\rangle}_{{U}}$) to those of the target V{V}, defined over a potentially different party set $\mathcal{P}^{\prime}$ (i.e., ${\langle{D}\rangle}_{{V}}$). The security guarantees of the Reshare algorithm are inherently tied to the specific instantiation of $\mathcal{F}_{\textnormal{{Reshare}}}$, which in turn depends on the threat models associated with both U{U} and V{V}. In practice, these instantiations vary based on the deployment setting. Detailed constructions and examples for clusters of 2, 3, and 4 parties are discussed in §3.6.
3.3 Threat Model
Given the hybrid nature of our framework, we adopt a mixed adversarial corruption model [BJMS20, HM20] in WW-FL, assuming a centralized adversary $\mathcal{A}$ that orchestrates the corruption [^3] This is a stronger notion than having multiple independent adversaries.. This setup allows the nature of corruption to vary across different MPC clusters, including the global server cluster. We formalize this via a corruption strategy $\mathsf{St}_{\delta}$ defined over each entity set $\delta\in\{\mathcal{G},\mathcal{M}_{i},\mathcal{C}_{i}\}_{i\in\mathsf{N}_{\scriptstyle\mathcal{M}}}$.
For instance, $\mathcal{A}$ may choose to maliciously corrupt certain MPC clusters $\mathcal{M}_{i}$ (reflecting an unjustified trust assumption in the cluster), while corrupting others, including the global servers $\mathcal{G}$, in a semi-honest manner. Within each entity set, $\mathcal{A}$ may corrupt a minority or a majority of parties, depending on the number of honest parties assumed (with the minimal assumption that each set contains at least one honest entity).
To reflect the hybrid structure of WW-FL composed of multiple MPC clusters, we constrain the adversary $\mathcal{A}$ to employ a uniform corruption type within each cluster – either semi-honest or malicious – but not a mixture of both. Allowing mixed corruption types within a single cluster [Can00, HMZ08], while a natural extension, is left as future work.
@@ 3.3 Threat Model
Table 4: Possible configurations of WW-FL in terms of security against various corruption strategies of the adversary. Here $0<t<\mathsf{N}_{\mathcal{M}}$.
@@
I-A semi-honest $\mathsf{N}_{\mathcal{M}}$ semi-honest semi-honest
I-B semi-honest $t$ malicious semi-honest
I-C semi-honest $\mathsf{N}_{\mathcal{M}}$ malicious semi-honest
II-A malicious $\mathsf{N}_{\mathcal{M}}$ semi-honest semi-honest
II-B malicious $t$ malicious semi-honest
II-C malicious $\mathsf{N}_{\mathcal{M}}$ malicious semi-honest
III-A semi-honest $\mathsf{N}_{\mathcal{M}}$ semi-honest malicious
III-B semi-honest $t$ malicious malicious
III-C semi-honest $\mathsf{N}_{\mathcal{M}}$ malicious malicious
IV-A malicious $\mathsf{N}_{\mathcal{M}}$ semi-honest malicious
IV-B malicious $t$ malicious malicious
IV-C malicious $\mathsf{N}_{\mathcal{M}}$ malicious malicious
The resulting security spectrum of WW-FL is summarized in Tab. 4, ranging from the weaker semi-honest cases to the strongest all-malicious setting, depending on the adversary’s corruption strategy across entities.
@@
We model the FL training process in WW-FL as an ideal functionality $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ which is depicted in Alg. 3. For simplicity, we assume that the clients have secret-shared their training data with their respective MPC clusters via the respective Share protocol. In each iteration $t$, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ receives the shares of the current model $W_{t-1}$ from the global servers ($\mathcal{G}$) and the shares of the training data from each cluster’s servers in $\mathcal{M}_{i}$ for $i\in\mathsf{N}_{\scriptstyle\mathcal{M}}$. Using these shares, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ reconstructs the current global model and the training data for each cluster. Next, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ trains the model $W_{t-1}$ using the data from each cluster, resulting in a cluster model. These cluster models are then combined using an aggregation algorithm to obtain the updated global model $W_{t}$ for this iteration. Finally, $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ generates the secret-shares of $W_{t}$ among the global servers ($\mathcal{G}$) and the cluster servers ($\mathcal{M}_{i}$) using their respective sharing semantics.
@@
0:  $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$, $\mathcal{G},\mathcal{M}$ $\mathcal{M}=\bigcup_{i}\mathcal{M}_{i}$; $i\in\mathsf{N}_{\scriptstyle\mathcal{M}}$
@@
0:  ${\langle W_{t}\rangle}_{\mathcal{G}}$, ${\langle W_{t}\rangle}_{\mathcal{M}_{i}}$$W_{t}$ – global model after iteration $t$
@@
3:   data reconstruction: $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ receives $\langle{D}_{t}^{\mathcal{M}_{i}}\rangle$} from $\mathcal{M}_{i}$ and reconstructs the cluster data ${D}_{t}^{\mathcal{M}_{i}}$${D}_{t}^{\mathcal{M}_{i}}$ corresponds to collective data at $\mathcal{M}_{i}$
@@
6:  aggregation: $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ aggregates $\{W_{t}^{\mathcal{M}_{i}}\}_{i\in\mathsf{N}_{\scriptstyle\mathcal{M}}}$ using the respective secure aggregation algorithm to obtain $W_{t}{}$.
@@
8:  secret-sharing (cluster servers): $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ generates ${\langle W_{t}\rangle}_{\mathcal{M}_{i}}$ from $W_{t}{}$ for $i\in\mathsf{N}_{\scriptstyle\mathcal{M}}$.
Security of WW-FL
We adopt the standard stand-alone model of [Can00], assuming a static adversary and synchronous communication over perfectly secure channels [^4] In practice, such communication channels can be instantiated via mutually authenticated TLS.. Our security analysis is carried out in a simulation-based framework [Lin17], which offers a modular and composable approach to proving security – particularly well-suited to the clustered architecture of the WW-FL framework.
Specifically, for most subprotocols, we first define the corresponding ideal functionalities. These functionalities formalize their expected behavior in an ideal execution with a trusted third party, as is standard in MPC. Security of the overall framework is then established via a hybrid proof: each subprotocol securely realizes its functionality, and the composition theorem allows us to combine these results without re-deriving the security of every underlying instantiation [Lin17]. Importantly, we do not assume the existence of corresponding “ideal implementations”. Instead, we instantiate the functionalities using existing MPC protocols with proven security guarantees, and compose them within WW-FL to implement more complex tasks.
Let $\mathcal{F}_{\textsf{MPC}}^{\gamma}$ denote the set of MPC ideal functionalities associated with the entity set $\gamma\in\{\mathcal{G},\,\mathcal{M}_{i}\}_{i\in\mathsf{N}_{\scriptstyle\mathcal{M}}}$. These functionalities vary depending on the role of the entity in the WW-FL framework. Specifically, for a standard FL setup, [^5] In the presence of defense mechanisms, $\mathcal{F}_{\textsf{MPC}}^{\gamma}$ may include additional functionalities as listed in Tab. 3; these are straightforward generalizations.
$\mathcal{F}_{\textsf{MPC}}^{\gamma}=\begin{cases}\{\mathcal{F}_{\textnormal{{Share}}}^{\mathcal{M}_{i}},\,\mathcal{F}_{\textnormal{{Reshare}}}^{\mathcal{M}_{i}},\,\mathcal{F}_{\textnormal{{Sample}}}^{\mathcal{M}_{i}},\,\mathcal{F}_{\textnormal{{Train}}}^{\mathcal{M}_{i}}\}&\text{if }\gamma=\mathcal{M}_{i},\\[5.0pt] \{\mathcal{F}_{\textnormal{{Share}}}^{\mathcal{G}},\,\mathcal{F}_{\textnormal{{Reshare}}}^{\mathcal{G}},\,\mathcal{F}_{\textnormal{{Agg}}}^{\mathcal{G}}\}&\text{if }\gamma=\mathcal{G}.\end{cases}$
Given these definitions, the overall security of the WW-FL framework is formally stated in Theorem 3.1.
@@
Let $\mathcal{F}_{\textsf{MPC}}^{\gamma}$ denote the set of ideal MPC functionalities for each entity set $\gamma\in\{\mathcal{G},\,\mathcal{M}_{i}\}_{i\in\mathsf{N}_{\scriptstyle\mathcal{M}}}$. Algorithm FL-Train (Alg. 1) securely realizes the $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ functionality (Alg. 3) in the $\mathcal{F}_{\textsf{MPC}}^{\gamma}$-hybrid model, assuming a static adversary $\mathcal{A}$ who corrupts the entities in $\delta\in\{\mathcal{G},\mathcal{M}_{i},\mathcal{C}_{i}\}_{i\in\mathsf{N}_{\scriptstyle\mathcal{M}}}$, according to a predefined corruption strategy $\mathsf{St}_{\delta}$.
Proof Sketch
The FL-Train algorithm (Alg. 1) begins with the generation of secret shares of the initial global model $W_{0}$ among the global servers $\mathcal{G}$. This operation is performed by the designated stakeholder responsible for deploying the WW-FL framework, using the Share algorithm instantiated for $\mathcal{G}$. The security of this sharing procedure is guaranteed by the ideal functionality $\mathcal{F}_{\textnormal{{Share}}}^{\mathcal{G}}$.
In an iteration $t\in[1,T]$, the global model $W_{t-1}$, that is secret-shared among the global servers must be securely redistributed to the servers within each Layer II cluster. This redistribution is carried out via the Reshare protocol executed by $\mathcal{G}$ with respect to each cluster $\mathcal{M}_{i}$, for all $i\in\mathsf{N}_{\scriptstyle\mathcal{M}}$. The security of this step follows from the ideal functionality $\mathcal{F}_{\textnormal{{Reshare}}}^{\mathcal{G}}$ (cf. Alg. 2), where the source party set is $\mathcal{P}=\mathcal{G}$, holding the secret-shared model ${\langle W_{t-1}\rangle}_{\mathcal{G}}$, and the target party set is $\mathcal{P}^{\prime}=\mathcal{M}_{i}$, which receives the reshared model ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}$.
In the next step of the FL-Train algorithm, the selected clients $\mathsf{C}_{i}^{j}$ associated with cluster $\mathcal{M}_{i}$ (i.e., those in the set $\mathcal{C}_{i}^{\sf U}$ securely sampled via the Sample algorithm) generate secret shares of their local data ${D}_{t}^{\mathsf{C}_{i}^{j}}$ among the servers in $\mathcal{M}_{i}$, denoted by ${\langle{D}_{t}^{\mathsf{C}_{i}^{j}}\rangle}_{\mathcal{M}_{i}}$. This sharing is performed using the Share algorithm instantiated for $\mathcal{M}_{i}$, and its security is ensured by the corresponding ideal functionality $\mathcal{F}_{\textnormal{{Share}}}^{\mathcal{M}_{i}}$. Subsequently, each cluster $\mathcal{M}_{i}$ performs ML training on its collective dataset ${\langle{D}_{t}\rangle}_{\mathcal{M}_{i}}$ using the Train algorithm, whose security is guaranteed by the corresponding ideal functionality $\mathcal{F}_{\textnormal{{Train}}}^{\mathcal{M}_{i}}$.
Even in the presence of malicious corruption, the ideal functionality $\mathcal{F}_{\textnormal{{Share}}}^{\mathcal{M}_{i}}$ guarantees that a malicious client cannot introduce inconsistent shares across the servers in $\mathcal{M}_{i}$. While a malicious client may tamper with its input data prior to the secret-sharing process, this behavior falls outside the scope of MPC itself and is typically addressed through complementary defense mechanisms (cf. §5 for defenses in WW-FL). As such, $\mathcal{F}_{\textnormal{{Train}}}^{\mathcal{M}_{i}}$ ensures the privacy and correctness of the ML training on data that has been secret-shared by clients. In cases where integrity violations arise due to adversarial behavior, the underlying secure aggregation protocol Agg, which realizes the ideal functionality $\mathcal{F}_{\textnormal{{Agg}}}^{\mathcal{G}}$, plays a critical role in detecting and mitigating such attacks.
In the next step, the local model ${\langle W_{t}^{i}\rangle}_{\mathcal{M}_{i}}$ is securely reshared with the global servers $\mathcal{G}$ using the Reshare protocol. The security of this resharing step is ensured by the ideal functionality $\mathcal{F}_{\textnormal{{Reshare}}}^{\mathcal{M}_{i}}$, where the source party set is $\mathcal{P}=\mathcal{M}_{i}$, holding the secret-shared cluster model ${\langle W_{t}^{i}\rangle}_{\mathcal{M}_{i}}$, and the target party set is $\mathcal{P}^{\prime}=\mathcal{G}$, which receives the reshared model ${\langle W_{t}^{i}\rangle}_{\mathcal{G}}$.
Finally, the global servers $\mathcal{G}$ execute the Agg algorithm to securely aggregate the models received from all clusters, i.e., $\{{\langle W_{t}^{i}\rangle}_{\mathcal{G}}\}_{i\in\mathsf{N}_{\scriptstyle\mathcal{M}}}$, and compute the updated global model in secret-shared form, denoted by ${\langle W_{t}\rangle}_{\mathcal{G}}$. The security of this aggregation step is guaranteed by the ideal functionality $\mathcal{F}_{\textnormal{{Agg}}}^{\mathcal{G}}$.
Therefore, assuming that all underlying functionalities are securely instantiated in accordance with the specified security assumptions for each set $\delta~\in~\{\mathcal{G},\mathcal{M}_{i},\mathcal{C}_{i}\}_{i\in\mathsf{N}_{\scriptstyle\mathcal{M}}}$ in WW-FL, the FL-Train algorithm (Alg. 1) securely realizes the ideal functionality $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ (Alg. 3). $\square$
3.4 Private Inference in WW-FL